# RIFT: Closed-Loop RL Fine-Tuning for Realistic and Controllable Traffic Simulation

### v2

@@ 4.1 Trajectory Generation for CBVs
4.2 Closed-Loop RL Fine-Tuning
While effective in capturing trajectory-level realism and route-level controllability, the open-loop IL pre-training approach suffers from covariate shift between training and deployment, resulting in error accumulation and unrealistic long-term behavior in closed-loop settings. Prior works have addressed this issue through RL PPO ; sutton1999reinforce or hybrid IL–RL fine-tuning zhang2023rtr ; peng2024improving , but typically optimize only the executed trajectory, ignoring other candidates and thereby risking mode collapse.
@@ 4.2 Closed-Loop RL Fine-Tuning
To preserve multimodality during closed-loop fine-tuning, we adopt GRPO-style group-relative optimization shao2024grpo , which explicitly evaluates all candidate modalities and assigns greater relative advantages to those that better align with user-preferred styles. This approach enhances style-level controllability while maintaining multimodality. Considering closed-loop dynamics, we evaluate simulated rollouts rather than raw candidate trajectories to mitigate plan–rollout deviation. Specifically, given $G=N_{\mathrm{ref}}\times N_{\mathrm{lon}}$ candidate trajectories $\mathcal{T}=\{\tau_{i}\}_{i=1}^{G}$ for a particular CBV, we perform short-horizon forward simulation dauner2023tuplan_garage to obtain rollouts $\widetilde{\mathcal{T}}=\{\widetilde{\tau}_{i}\}_{i=1}^{G}$, considered jointly as a group. Each rollout is evaluated by a user-defined state-wise reward model $\mathrm{StateWiseRM}$, from which we compute the corresponding discounted returns $\mathcal{R}=\{r_{i}\}_{i=1}^{G}$ and derive the group-relative advantages $\mathcal{A}=\{\hat{A}_{i}\}_{i=1}^{G}$ as follows:
@@
$r_{i}=\sum_{t=0}^{T}\gamma^{t}\left[\mathrm{StateWiseRM}(\widetilde{\tau}_{i}^{t})\right],\ \hat{A}_{i}=\frac{r_{i}-\mathrm{mean}(\mathcal{R})}{\mathrm{std}(\mathcal{R})},\quad\forall i=1,\dots,G$ (6)
Here, $\hat{A}_{i}$ measures each candidate rollout’s performance relative to the group, promoting high-return rollouts while preserving multimodality. Given the group-relative advantage $\hat{A}_{i}$, we derive the standard GRPO objective in our framework as:
@@
$\begin{split}\mathcal{J}_{\mathrm{GRPO}}(\theta)&=\mathbb{E}_{[\{\tau_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}(s)}]}\\ &\frac{1}{G}\sum_{i=1}^{G}\left\{\min\left[\frac{\pi_{\theta}(\tau_{i}|s)}{\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)}\hat{A}_{i},\mathrm{clip}\left(\frac{\pi_{\theta}(\tau_{i}|s)}{\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)},1-\epsilon,1+\epsilon\right)\hat{A}_{i}\right]-\beta\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right]\right\},\end{split}$ (7)
where ${\pi}_{\theta}$ is the current planning model, ${\pi}_{\mathrm{ref}}$ is the reference model, $\epsilon$ and $\beta$ are hyper-parameters, and $\rho_{i}(\theta)=\pi_{\theta}(\tau_{i}|s)/{\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)}$ is the trajectory-level importance sampling ratio.
Due to covariate shift in closed-loop deployment, the trajectory scoring head—lacking reference-line priors—is more susceptible to distribution mismatch than the trajectory generation head. As a result, the pre-trained scoring head may produce unstable or misaligned results. In contrast, the generation head, anchored by route-level priors, remains robust, consistently generating realistic and multimodal trajectories. To preserve this capability while enhancing style-level controllability and mitigating covariate shift, we freeze the generation head and fine-tune only the scoring head toward user-preferred styles. In this setting, constraining the scoring head with KL regularization to the IL pre-trained model would anchor learning to a biased reference, thereby hindering adaptation. We therefore remove the KL term, allowing the scoring head to adapt freely toward user preferences while retaining the trajectory-level realism and route-level controllability established during IL pre-training. Moreover, closed-loop RL fine-tuning enables the model to recover from non-expert states encountered during deployment, further alleviating covariate shift in trajectory scoring.
Removing KL regularization can introduce instability during fine-tuning. PPO-Clip PPO mitigates this by constraining policy updates via a clipped importance sampling ratio, effectively restricting shifts to a trust region. However, this mechanism is insufficient in our group-relative setting: the original clipping targets only the executed action in policy learning, whereas we evaluate an entire set of trajectory candidates. Here, the trajectory-level importance ratio $\rho_{i}(\theta)=\pi_{\theta}(\tau_{i}|s)/{\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)}$ can grow arbitrarily large, when $\pi_{\theta}(\tau_{i}|s)$ dominates $\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)$ and $\hat{A}_{i}<0$, the product $\rho_{i}(\theta)\hat{A}_{i}$ may yield large negative values, causing high variance and unstable updates. To address this while maintaining alignment with user-preferred styles, we adopt the dual-clip mechanism from Dual-Clip PPO ye2020dual_clip ; gao2021learning , where a constant $c>1$ bounds clipped negative advantages from below. Building on the theoretical analysis in Appendix˜A, we define the resulting RIFT objective as:
@@
$\begin{split}\mathcal{J}_{\mathrm{RIFT}}(\theta)&=\mathbb{E}_{[\{\tau_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}(s)}]}\\ &\frac{1}{G}\sum_{i=1}^{G}\left\{\max\left[\min\left[\frac{\pi_{\theta}(\tau_{i}|s)}{\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)}\hat{A}_{i},\mathrm{clip}\left(\frac{\pi_{\theta}(\tau_{i}|s)}{\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)},1-\epsilon,1+\epsilon\right)\hat{A}_{i}\right],c\hat{A}_{i}\right]\right\}.\end{split}$ (8)
@@ 5 Experiment
This section systematically addresses the following research questions: Q1: How does RIFT compare with representative baselines in terms of the realism and controllability of the generated traffic scenarios? Q2: How can the generated traffic scenario be effectively utilized to support downstream autonomous driving tasks? Q3: What are the contributions of each component within RIFT to the overall performance, and how well is style-level controllability preserved under different user-specified driving styles? We first describe the experimental setup, including the simulation environment, baseline methods, and evaluation metrics, followed by a detailed analysis of the results in response to the proposed research questions.
@@ 5.1 Experiment Setups
Under the dual-stage AV-centric simulation framework, we adopt Pluto cheng2024pluto as our planning model for its well-established performance and open-source implementation. To ensure fair comparison, we use the official IL pre-trained checkpoint provided by Pluto, trained on the nuPlan dataset caesar2021nuplan . Simulations are conducted in CARLA carla , leveraging Bench2Drive jia2024bench2drive to support AV-centric closed-loop simulation and evaluation. Implementation details, training protocols, and evaluation settings are described in Appendix˜B.
Baseline. To systematically evaluate the effectiveness of RIFT in closed-loop fine-tuning, we compare it against the following baselines, with implementation details provided in Section˜B.5.
All methods are fine-tuned on the same module—the scoring head—to ensure fair comparisons, while isolating style-level controllability from trajectory-level realism and route-level controllability, as confirmed by the ablation studies in Section˜5.4. Following the realism standards of the Sim Agent Challenge in WOSAC montali2023wosac , we adopt a normal style reward for all RL-based baselines, with details in Section˜B.7. Results under an aggressive style reward are also reported in Section˜5.4.
Metrics. Building on the WOSAC evaluation framework montali2023wosac , we categorize our evaluation metrics into three groups: kinematic metrics, interaction metrics, and map metrics. Kinematic metrics focus on distributional motion properties (S-SW, S-WD, A-SW), as inspired by chen2024review ; huang2025IntSim , with the absence of ground-truth trajectories in CARLA precluding displacement-based measures (e.g., ADE, FDE). Interaction metrics evaluate agent interactions through collision frequency (Collision Per Kilometer, CPK), driving efficiency (Route Progress, RP), and safety-critical measures, including 2D-TTC guo20232dttc and ACT venthuruthiyil2022act . Map metrics evaluate adherence to road geometry through the Off-Road Rate (ORR). Collectively, these metrics provide a comprehensive evaluation of realism and controllability in closed-loop simulation; detailed definitions are provided in Section˜B.8.
Table 1: Comparison in Controllability and Realism. Metrics are evaluated under the PDM-Lite Beibwenger2024pdmLite AV setting across three random seeds, with the best and the second-best results highlighted accordingly.
@@
Pluto cheng2024pluto IL 0.88 $\pm$ 0.01 5.81 $\pm$ 0.06 0.90 $\pm$ 0.01 5.06 $\pm$ 2.69 564.14 $\pm$ 114.41 2.50 $\pm$ 1.48 2.44 $\pm$ 1.39 0.24 $\pm$ 0.15
PPO chen2024frea RL 0.95 $\pm$ 0.01 4.45 $\pm$ 0.15 0.89 $\pm$ 0.02 13.95 $\pm$ 2.34 409.51 $\pm$ 30.38 2.59 $\pm$ 1.60 2.52 $\pm$ 1.57 9.17 $\pm$ 2.39
FREA chen2024frea RL 0.93 $\pm$ 0.01 5.10 $\pm$ 0.14 0.93 $\pm$ 0.01 30.42 $\pm$ 5.28 292.81 $\pm$ 68.54 2.71 $\pm$ 1.40 2.67 $\pm$ 1.41 9.01 $\pm$ 2.09
FPPO-RS chen2024frea RL 0.87 $\pm$ 0.01 5.80 $\pm$ 0.11 0.80 $\pm$ 0.03 21.39 $\pm$ 3.23 356.79 $\pm$ 26.19 2.55 $\pm$ 1.69 2.53 $\pm$ 1.68 8.60 $\pm$ 0.25
@@
RS-Pluto peng2024improving SFT+RLFT 0.93 $\pm$ 0.00 5.40 $\pm$ 0.15 0.92 $\pm$ 0.01 4.11 $\pm$ 3.90 819.40 $\pm$ 74.07 2.27 $\pm$ 1.45 2.23 $\pm$ 1.43 1.05 $\pm$ 0.31
RTR-Pluto zhang2023rtr SFT+RLFT 0.85 $\pm$ 0.00 6.24 $\pm$ 0.16 0.81 $\pm$ 0.03 6.98 $\pm$ 2.59 481.60 $\pm$ 70.19 2.55 $\pm$ 1.60 2.47 $\pm$ 1.51 0.08 $\pm$ 0.09
@@
Figure 3: Speed and Acceleration Distribution. RL-based methods produce interactive but often unnatural behaviors, while supervised methods typically lead to conservative behavior. In contrast, RIFT achieves higher interactivity, reflected by increased average speed and acceleration, and more realistic behavior from the distributional perspective, avoiding unnecessary hesitation while keeping safe interactions.
@@ 5.2 Realistic and Controllable Traffic Scenario Generation (Q1)
Main Results. To address Q1, we evaluate the controllability and realism of the generated scenario across different CBV methods, with results summarized in Table˜1. RIFT consistently outperforms all baselines in both aspects across most settings. While supervised learning methods achieve slightly lower CPK and ORR, this improvement is primarily due to their inherently conservative behavior, derived from the expert PDM-Lite Beibwenger2024pdmLite , which prioritizes safety by avoiding risky maneuvers.
@@ 5.3 Generated Traffic Scenarios for Closed-Loop AV Evaluation (Q2)
To address Q2, we assess the suitability of traffic scenarios generated by different CBV methods for closed-loop AV evaluation. Following KING king , we adopt PDM-Lite Beibwenger2024pdmLite —a rule-based planner with privileged access—as a reference to evaluate two key scenario properties: feasibility, measured by Driving Score (DS), and naturalness, captured by our proposed Blocked Rate (BR). A high DS indicates that the AV can reliably complete the scenario, while a low BR reflects realistic interactions without excessive obstruction from surrounding vehicles. Together, DS and BR offer a principled basis for evaluating scenario quality.
To further assess each scenario’s ability to reveal weaknesses in learning-based planners, we compare the performance of PlanT Renz2022PlanT , UniAD hu2023_uniad , and VAD jiang2023vad against PDM-Lite. As these models are more sensitive to subtle or adversarial interactions, informative scenarios should induce noticeable performance drops. As shown in Table˜3, traffic generated by RIFT achieves the highest DS and lowest BR under PDM-Lite, while also causing the most significant degradation in all learning-based planners. These results demonstrate that RIFT produces realistic and well-structured scenarios that are effective at exposing the limitations of modern AV systems. See Appendix˜C for further details.
@@
Method PDM-Lite Beibwenger2024pdmLite PlanT Renz2022PlanT UniAD hu2023_uniad VAD jiang2023vad
@@
Pluto cheng2024pluto 77.84 $\pm$ 2.20 23.33 $\pm$ 5.77 42.52 $\pm$ 4.72 -35.32 73.73 $\pm$ 1.24 -4.11 66.87 $\pm$ 2.11 -10.97
PPOchen2024frea 76.26 $\pm$ 0.12 30.00 $\pm$ 0.00 36.39 $\pm$ 1.11 -39.87 69.79 $\pm$ 1.41 -6.47 67.64 $\pm$ 1.27 -8.62
FREAchen2024frea 83.53 $\pm$ 0.13 20.00 $\pm$ 0.00 39.61 $\pm$ 1.34 -43.92 69.29 $\pm$ 5.22 -14.24 67.57 $\pm$ 5.37 -15.96
FPPO-RSchen2024frea 83.52 $\pm$ 0.09 20.00 $\pm$ 0.00 38.85 $\pm$ 4.91 -44.67 75.13 $\pm$ 5.18 -8.39 69.15 $\pm$ 2.79 -14.37
@@
RS-Pluto peng2024improving 89.32 $\pm$ 1.41 13.33 $\pm$ 5.77 42.05 $\pm$ 4.08 -47.27 80.62 $\pm$ 0.78 -8.70 69.48 $\pm$ 5.02 -19.84
RTR-Pluto zhang2023rtr 87.64 $\pm$ 1.56 10.00 $\pm$ 0.00 40.08 $\pm$ 2.38 -47.56 77.69 $\pm$ 2.82 -9.95 66.27 $\pm$ 4.53 -21.37
@@
Table 3: Ablation Study on RIFT. Evaluation under PDM-Lite Beibwenger2024pdmLite AV setting with three random seeds.

### v3

@@ 4.1 Route-Level Interaction Analysis
4.2 Group-Relative RL Fine-Tuning
Open-loop IL pre-training offers trajectory-level realism and route-level controllability; however, it inevitably suffers from covariate shift in closed-loop deployment, causing error accumulation and unrealistic long-term behaviors. Existing RL [60] and hybrid IL–RL methods [26] partially mitigate covariate shift, but their optimization is restricted to the executed rollout, disregarding alternative candidates and degrading multimodality. More critically, covariate shift induces asymmetric degradation across model components: under the generation–selection paradigm, the generation head, conditioned on route-level priors, remains robust and consistently produces realistic multimodal candidates, whereas the scoring head, trained solely through imitation, is more vulnerable to distribution mismatch. These challenges motivate three key requirements for fine-tuning: (i) preserving multimodality, (ii) addressing asymmetric covariate shift, and (iii) ensuring stable policy improvement. We address these requirements through a unified framework that combines group-relative optimization, asymmetry-aware fine-tuning, and dual-clip stabilization.
@@ 4.2 Group-Relative RL Fine-Tuning
To preserve multimodality, we adopt group-relative formulation [24], which evaluates all candidate modalities within the group and assigns higher relative advantages to those better aligned with user-preferred styles. Considering closed-loop dynamics, we evaluate simulated rollouts rather than raw trajectories to mitigate plan–rollout deviation. Specifically, given $\smash{G=N_{\mathrm{ref}}\times N_{\mathrm{lon}}}$ candidate trajectories $\smash{\mathcal{T}=\{\tau_{i}\}_{i=1}^{G}}$ for a CBV at state $s$, we conduct forward simulation [61] (see Appendix B.6) to obtain rollouts $\smash{\widetilde{\mathcal{T}}=\{\widetilde{\tau}_{i}\}_{i=1}^{G}}$. Each rollout is evaluated by a user-defined state-wise reward model $\mathrm{StateWiseRM}$, yielding the corresponding discounted returns $\smash{\mathcal{R}=\{R_{i}\}_{i=1}^{G}}$ from which we derive the group-relative advantages $\smash{\mathcal{A}=\{\hat{A}_{i}\}_{i=1}^{G}}$ as follows:
@@
$R_{i}(s)=\sum_{t=0}^{T}\gamma^{t}\left[\mathrm{StateWiseRM}(\widetilde{\tau}_{i}^{t},s)\right],\ \qquad\hat{A}_{i}(s)=\frac{R_{i}(s)-\mathrm{mean}(\mathcal{R})}{\sqrt{\mathrm{Var}(\mathcal{R})+\varepsilon}}.$ (6)
Here, $\hat{A}_{i}$ quantifies the performance of each rollout relative to the group, promoting high-return rollouts without suppressing alternative modes.
In standard GRPO [24], sampling from the old policy implicitly induces old-policy weighting. Extending this to our enumerated setting involves averaging terms weighted by $\pi_{\theta_{\mathrm{old}}}$ in conjunction with the importance ratio $\smash{\rho_{i}(\theta)=\pi_{\theta}(\tau_{i}\!\mid\!s)/\pi_{\theta_{\mathrm{old}}}(\tau_{i}\!\mid\!s)}$, which yields a low-variance estimate of the old-policy expectation over the enumerated support. The aggregated objective is:
@@
$\begin{split}\mathcal{J}(\theta)=\mathbb{E}_{s\sim\mathcal{D}}\Bigg[\sum_{i=1}^{G}\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)\min\left[\rho_{i}(\theta)\hat{A}_{i},\mathrm{clip}\left(\rho_{i}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{i}\right]\Bigg]-\beta\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right],\end{split}$ (7)
where $\smash{\pi_{\mathrm{ref}}}$ denotes the IL pre-trained model. While exact over the enumerated support, this scheme overemphasizes frequent modes and under-represents rare but high-return ones, causing mode collapse and reduced diversity. To balance modality contributions, we adopt an equal-weight objective:
@@
$\begin{split}\mathcal{J}(\theta)=\mathbb{E}_{s\sim\mathcal{D}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\min\left[\rho_{i}(\theta)\hat{A}_{i},\mathrm{clip}\left(\rho_{i}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{i}\right]\Bigg]-\beta\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right].\end{split}$ (8)
Under equal weighting, $\rho_{i}(\theta)$ regulates candidate updates rather than serving as a pure importance weight, removing old-policy bias and yielding balanced updates that preserve multimodality.
To address asymmetric covariate shift, we freeze the generation head to retain trajectory-level realism and fine-tune only the scoring head to enhance style-level controllability. In this setting, constraining the scoring head with the KL term to the IL pre-trained model would anchor learning to a biased reference, thereby hindering adaptation. We therefore remove the KL term, allowing the scoring head to adapt freely while leveraging the stable candidates provided by the frozen generation head.
Removing the KL term improves flexibility but raises stability concerns. Although the clipped-ratio mechanism in PPO constrains update magnitude, it proves insufficient in the group-relative setting. Specifically, when a rare trajectory under the old policy receives a higher probability from the current policy despite a negative advantage, the product $\smash{\rho_{i}(\theta)\hat{A}_{i}}$ can become disproportionately large and destabilize learning. To address this, we incorporate the dual-clip surrogate from Dual-Clip PPO [62, 63], which lower-bounds clipped negative advantages. This establishes a trust-region-like constraint that guarantees bounded per-candidate updates (see Lemma˜A.3), thereby preventing extreme negative shifts while preserving responsiveness to user-preferred styles. The resulting surrogate objective, termed RIFT, is
$\mathcal{J}_{\mathrm{RIFT}}(\theta)$ $=\mathbb{E}_{s\sim\mathcal{D}}\!\left[\frac{1}{G}\sum_{i=1}^{G}\;\psi\big(\rho_{i}(\theta),\,\hat{A}_{i}\big)\right],$ (9)
[…]
This objective integrates multimodality preservation, asymmetry-aware fine-tuning, and stable optimization into a unified framework, enhancing style-level controllability and mitigating covariate shift while retaining trajectory-level realism and route-level controllability (analysis in Appendix˜A).
@@ 5 Experiment
This section systematically addresses the following research questions: Q1: How does RIFT compare with representative baselines in terms of the realism and controllability of the generated traffic scenarios? Q2: How can the generated traffic scenario be effectively utilized to support downstream autonomous driving tasks? Q3: How do the components of RIFT contribute to overall performance, and to what extent is style-level controllability preserved under varying user-specified driving styles?
@@ 5.1 Experiment Setups
Under the dual-stage AV-centric simulation framework, we adopt Pluto [56] as our planning model for its well-established performance and open-source implementation. To ensure fair comparison, we use the official IL pre-trained checkpoint provided by Pluto, trained on the nuPlan dataset [15]. Simulations are conducted in CARLA [64], leveraging Bench2Drive [49] to support AV-centric closed-loop simulation and evaluation. Implementation details, training protocols, and evaluation settings are described in Appendix˜B.
Table 1: Comparison in Controllability and Realism. Metrics are evaluated under the PDM-Lite [65] AV setting across three random seeds, with the best and the second-best results highlighted accordingly.
@@
Pluto IL 0.88 $\pm$ 0.01 5.81 $\pm$ 0.06 0.90 $\pm$ 0.01 5.06 $\pm$ 2.69 564.14 $\pm$ 114.41 2.50 $\pm$ 1.48 2.44 $\pm$ 1.39 0.24 $\pm$ 0.15
PPO RL 0.95 $\pm$ 0.01 4.45 $\pm$ 0.15 0.89 $\pm$ 0.02 13.95 $\pm$ 2.34 409.51 $\pm$ 30.38 2.59 $\pm$ 1.60 2.52 $\pm$ 1.57 9.17 $\pm$ 2.39
FREA RL 0.93 $\pm$ 0.01 5.10 $\pm$ 0.14 0.93 $\pm$ 0.01 30.42 $\pm$ 5.28 292.81 $\pm$ 68.54 2.71 $\pm$ 1.40 2.67 $\pm$ 1.41 9.01 $\pm$ 2.09
FPPO-RS RL 0.87 $\pm$ 0.01 5.80 $\pm$ 0.11 0.80 $\pm$ 0.03 21.39 $\pm$ 3.23 356.79 $\pm$ 26.19 2.55 $\pm$ 1.69 2.53 $\pm$ 1.68 8.60 $\pm$ 0.25
@@
RS-Pluto SFT+RLFT 0.93 $\pm$ 0.00 5.40 $\pm$ 0.15 0.92 $\pm$ 0.01 4.11 $\pm$ 3.90 819.40 $\pm$ 74.07 2.27 $\pm$ 1.45 2.23 $\pm$ 1.43 1.05 $\pm$ 0.31
RTR-Pluto SFT+RLFT 0.85 $\pm$ 0.00 6.24 $\pm$ 0.16 0.81 $\pm$ 0.03 6.98 $\pm$ 2.59 481.60 $\pm$ 70.19 2.55 $\pm$ 1.60 2.47 $\pm$ 1.51 0.08 $\pm$ 0.09
@@
Figure 3: Speed and Acceleration Distribution. RL-based methods tend to be interactive but unnatural, whereas supervised methods are overly conservative. RIFT strikes a balance, yielding higher interactivity with realistic distributional profiles, reducing hesitation while maintaining safe interactions.
Baseline. To systematically evaluate the effectiveness of RIFT in traffic simulation, we compare it against the following baselines, with implementation details provided in Section˜B.5.
  • Pure RL/IL: Methods trained solely with RL or IL, without fine-tuning, including Pluto [56], as well as FREA, FPPO-RS, and PPO, all from [18].
  • RLFT/SFT: Methods that fine-tune the pre-trained Pluto model using either RL or supervised objectives, including PPO-Pluto [60], REINFORCE-Pluto [66], GRPO-Pluto [24], and SFT-Pluto.
  • Hybrid: Methods that combine RL and supervised fine-tuning, including RTR-Pluto [25] and RS-Pluto [26].
All methods are fine-tuned on the scoring head to ensure fair comparisons, while isolating style-level controllability from trajectory-level realism and route-level controllability, as confirmed by the ablation studies in Section˜5.4. Following the realism standards of the Sim Agent Challenge in WOSAC [42], we adopt a normal style reward for all RL-based baselines, with details in Section˜B.7. Results under an aggressive style reward are reported in Section˜5.4.
Metrics. Building on the WOSAC evaluation framework, we categorize our evaluation metrics into three groups: kinematic metrics, interaction metrics, and map metrics. Kinematic metrics capture distributional motion properties (S-SW, S-WD, A-SW), as in [67], with the absence of ground-truth trajectories in CARLA precluding displacement-based measures (e.g., ADE, FDE). Interaction metrics evaluate agent interactions through collision frequency (Collision Per Kilometer, CPK), driving efficiency (Route Progress, RP), and safety-critical measures, including 2D-TTC [68] and ACT [69]. Map metrics evaluate adherence to road geometry through the Off-Road Rate (ORR). Collectively, these metrics comprehensively evaluate realism and controllability in closed-loop simulation; detailed definitions are in Section˜B.8.
@@ 5.2 Realistic and Controllable Traffic Scenario Generation (Q1)
Main Results. To address Q1, we evaluate the controllability and realism of the generated scenario across CBV methods, with results summarized in Table˜1. RIFT consistently outperforms all baselines in both aspects across most settings. While supervised learning methods achieve slightly lower CPK and ORR, this improvement is primarily due to their inherently conservative behavior, derived from the expert PDM-Lite [65], which prioritizes safety by avoiding risky maneuvers.
@@ 5.3 Generated Traffic Scenarios for Closed-Loop AV Evaluation (Q2)
To address Q2, we assess the suitability of traffic scenarios generated by different CBV methods for closed-loop AV evaluation. Following KING [17], we adopt PDM-Lite [65]—a rule-based planner with privileged access—as a reference to evaluate two key scenario properties: feasibility, measured by Driving Score (DS), and naturalness, captured by our proposed Blocked Rate (BR). A high DS indicates that the AV can reliably complete the scenario, while a low BR reflects realistic interactions without excessive obstruction from surrounding vehicles. Together, DS and BR offer a principled basis for evaluating scenario quality.
To further assess the ability of each scenario to reveal weaknesses in learning-based planners, we compare PlanT [70], UniAD [53], and VAD [54] with PDM-Lite. As these models are sensitive to subtle or adversarial interactions, informative scenarios should induce noticeable performance drops. As shown in Table˜3, traffic generated by RIFT achieves the highest DS and lowest BR under PDM-Lite, while also causing the largest degradation across all learning-based planners. These results confirm that RIFT generates interactive and feasible scenarios that effectively expose limitations of modern AV systems. See Appendix˜C for detailed results.
@@
Method PDM-Lite PlanT UniAD VAD
@@
Pluto 77.84 $\pm$ 2.20 23.33 $\pm$ 5.77 42.52 $\pm$ 4.72 -35.32 73.73 $\pm$ 1.24 -4.11 66.87 $\pm$ 2.11 -10.97
PPO 76.26 $\pm$ 0.12 30.00 $\pm$ 0.00 36.39 $\pm$ 1.11 -39.87 69.79 $\pm$ 1.41 -6.47 67.64 $\pm$ 1.27 -8.62
FREA 83.53 $\pm$ 0.13 20.00 $\pm$ 0.00 39.61 $\pm$ 1.34 -43.92 69.29 $\pm$ 5.22 -14.24 67.57 $\pm$ 5.37 -15.96
FPPO-RS 83.52 $\pm$ 0.09 20.00 $\pm$ 0.00 38.85 $\pm$ 4.91 -44.67 75.13 $\pm$ 5.18 -8.39 69.15 $\pm$ 2.79 -14.37
@@
RS-Pluto 89.32 $\pm$ 1.41 13.33 $\pm$ 5.77 42.05 $\pm$ 4.08 -47.27 80.62 $\pm$ 0.78 -8.70 69.48 $\pm$ 5.02 -19.84
RTR-Pluto 87.64 $\pm$ 1.56 10.00 $\pm$ 0.00 40.08 $\pm$ 2.38 -47.56 77.69 $\pm$ 2.82 -9.95 66.27 $\pm$ 4.53 -21.37
@@
Table 3: Ablation Study on RIFT. Evaluation under PDM-Lite AV setting with three random seeds.